# Constraints on Axion-Like Particles from 16.5 Years of Fermi-LAT Data and Prospects for VLAST

### v1

@@ Abstract
Axion-like particles (ALPs), hypothetical particles beyond the Standard Model, are considered as promising dark matter candidates. ALPs can convert into photons and vice versa in a magnetic field via the Primakoff effect, potentially generating detectable oscillation in $\gamma$-ray spectra. This study analyzes 16.5 years of data from the Fermi Large Area Telescope (Fermi-LAT) on NGC 1275, the brightest galaxy in the Perseus cluster, to constrain the ALP parameter space. Our results improve the previous 95% exclusion limits of the photon-ALP coupling $g_{a\gamma}$ by a factor of 2 in the ALP mass range of $4\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$. Moreover, we investigate the projected sensitivity of the future Very Large Area $\gamma$-ray Space Telescope (VLAST) on searching for ALPs. We find that (i) the expected sensitivity on the ALP-photon coupling can be stronger than that from the upcoming International Axion Observatory (IAXO) in the ALP mass range of $2\times 10^{-11}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-7}\,\mathrm{eV}$, with the best sensitivity of $g_{a\gamma}\sim 7\times 10^{-13}\,\mathrm{GeV^{-1}}$ at $m_{a}\sim 2\times 10^{-10}\,\mathrm{eV}$; (ii) VLAST can extend the sensitivity of the ALP masses below $5\times 10^{-12}\,\mathrm{eV}$, where the ALP-photon coupling $g_{a\gamma}\gtrsim 1.5\times 10^{-11}\,\mathrm{GeV^{-1}}$ will be excluded; (iii) the entire parameter space of ALP accounting for TeV transparency can be fully tested. These results demonstrate that VLAST will offer an excellent opportunity for ALPs searches.
@@ I Introduction
The strong CP problem in quantum chromodynamics (QCD) motivated the proposal of the axion, a hypothetical particle beyond the Standard Model (SM) [1, 2, 3, 4], with the mass postulated to be directly proportional to its coupling with photons. More general axion-like particles (ALPs) predicted in many string-theory-motivated extensions of the SM [5, 6, 7], share similarities with axions but are not necessarily tied to the strong CP problem, allowing independent variation of mass and coupling  [^1] For convenience, we use ALPs to refer to the axions and ALPs in the following context.. Besides, the ALPs have been proposed as one of the promising dark matter (DM) candidates, whose cosmological abundance can be naturally consistent with observational data via the non-thermal production mechanism [8, 9, 10, 11, 12, 13, 14]. Additionally, ALPs also offer a potential solution to various astrophysical anomalies, such as the apparent transparency of TeV photons from extragalactic sources [15, 16, 17, 18]. These motivate the great efforts to detect ALPs through ground-based experiments and astronomical observations  [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@
Ground-based experiments can be categorized into three main types. The first, light-shining-through-walls (LSW) experiments, such as the Any Light Particle Search (ALPS) at DESY [36, 37], aim to detect ALPs by shining light through a barrier and observing its conversion into axions. The second type focuses on axion dark matter detection, exemplified by the Axion Dark Matter eXperiment (ADMX) [38, 39, 40, 41, 42, 43, 44, 45], which uses a strong magnetic field to convert dark matter axions into detectable microwave photons. The third type involves solar axion detection, represented by the Axion Solar Telescope (CAST) at CERN [46] and the upcoming International Axion Observatory (IAXO) [47], which aims to greatly enhance sensitivity to solar axions.
Astronomical observations, especially with $\gamma$-ray telescopes, have been believed to be capable of advancing the search for ALPs and effectively narrowing down the viable parameter space for ALP [48, 49, 50]. Using six years of the Fermi Large Area Telescope (Fermi-LAT) observation on NGC 1275, the coupling above $5\times 10^{-12}\,\mathrm{GeV^{-1}}$ has been excluded for ALP mass $5\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$ [51]. Similarly, the $\gamma$-ray observation for SN 1987A rules out the coupling above $2\times 10^{-13}\,\mathrm{GeV^{-1}}$ for ALP mass $m_{a}\lesssim 1\times 10^{-9}\,\mathrm{eV}$ [50]. Further constraints come from the $\gamma$-ray spectra of extragalactic core-collapse supernovae and bright Fermi flat-spectrum radio quasars (3C 454.3, 3C 279, and CTA 102) [52, 53]. The High Energy Stereoscopic System (H.E.S.S.) collaboration, using observation from PKS 2155--304, excludes the coupling above $2.1\times 10^{-11}\,\mathrm{GeV^{-1}}$ for ALP mass $1.5\times 10^{-8}\,\mathrm{eV}\lesssim m_{a}\lesssim 6\times 10^{-8}\,\mathrm{eV}$ [54]. Notably, Ref. [55] uses eight years of Fermi-LAT observations on PKS 2155--304 to exclude a previously allowed “hole-like” parameter region in the Ref. [51]. Based on H.E.S.S. observations, Ref. [56] analyzes the $\gamma$-ray spectra of TeV sources in the Galactic plane, excluding certain parameter space regions and further providing projected sensitivity for the next-generation Cherenkov Telescope Array (CTA). Ref. [57] combines GeV (Fermi-LAT) and TeV (MAGIC, VERITAS, and H.E.S.S.) data from three bright supernova remnants to search for ALPs. Ref. [58] performs high-resolution polarimetric measurements of radiation near the Galactic Center’s supermassive black hole (SMBH) using a subset of the Event Horizon Telescope (EHT) array, excluding the coupling above $1\times 10^{-12}\,\mathrm{GeV^{-1}}$ for ALP mass $1\times 10^{-19}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-18}\,\mathrm{eV}$.
@@ II Photon-ALP Oscillation
Table 1: The fundamental model parameters, including those for the double $\beta$-model and the magnetic field model.
@@
Figure 1: The survival probability of photons from NGC 1275 under different ALP parameters based on one realization of Gaussian turbulent magnetic field, the “Dominguez” EBL model and the “Jansson12” GMF model. In the left panel, with the fixed coupling ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV^{-1}}$), the photon survival probability is presented for different ALP masses (red solid line: $m_{a}\ll\omega_{\mathrm{pl}}$; green solid line: $m_{a}=1\times 10^{-9}\,\mathrm{eV}$; blue solid line: $m_{a}=1\times 10^{-6}\,\mathrm{eV}$). When the ALP mass is significantly smaller than the plasma frequency, its effect on the photon survival probability becomes negligible. In the right panel, with the ALP mass fixed ($m_{a}=1\times 10^{-9}\,\mathrm{eV}$), the photon survival probability is presented for different couplings (red solid line: $g_{a\gamma}=1\times 10^{-13}\,\mathrm{GeV^{-1}}$; green solid line: $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$; blue solid line: $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV^{-1}}$).
Observational evidence indicates the ubiquitous presence of magnetic fields in Perseus clusters (with the typical strength of $B\sim\mathrm{10\,\mu G}$) and the Milky Way ($B\sim\mathrm{1\,\mu G}$). The $\gamma$-ray photons emitted from NGC 1275 propagate through these magnetic fields and undergo mutual conversions with ALPs via the Primakoff effect. This process, known as photon-ALP oscillation, can introduce irregular oscillatory features into the observed $\gamma$-ray energy spectra of the sources. Such features which are difficult to produce through known astrophysical processes, serve as a smoking-gun signature for the existence of ALPs. Assuming a region of size $\ell$ with a uniform magnetic field, the probability of a photon with energy $E_{\gamma}$ converting into an ALP is given by $P_{\gamma\to a}$ [64, 65, 66, 15, 48, 67, 68, 69, 70, 22, 71], and the photon survival probability $P\left(g_{a\gamma},m_{a},E_{\gamma}\right)$ is
@@
where $\omega_{\mathrm{pl}}$ is the plasma frequency:
@@
This conversion process exhibits strong energy dependence, as detailed in Figure 1:
  • $\mathbf{E_{\gamma}\ll E_{c}:}$ the photon-ALP oscillation is weak, with a low conversion probability, and the photon survival probability is close to 1.
  • $\mathbf{E_{\gamma}\sim E_{c}:}$ photons and ALPs undergo strong mutual conversion, leading to obvious oscillation.
  • $\mathbf{E_{\gamma}\gg E_{c}:}$ the conversion probability is still large, however, the photon survival probability in Equation (2) mildly varies with $E_{\gamma}$ and becomes gradual.
To distinguish from known astrophysical processes, we focus on searching for the oscillation features with $E_{\gamma}\sim E_{c}$. For typical astrophysical and ALP parameters ($B_{\mathrm{T}}=1\,\mu G,n_{e}=0.01\,\mathrm{cm^{-3}}$ for galaxy clusters or $n_{e}=0.1\,\mathrm{cm^{-3}}$ for Milky Way, $m_{a}=1\times 10^{-9}\,\mathrm{eV},g_{a\gamma}=1\times 10^{-10}\,\mathrm{GeV^{-1}}$), the resulting critical energy $E_{c}\approx 250\,\mathrm{MeV}$ falls within the optimal detectable energy ranges of both Fermi-LAT and VLAST, rendering them highly suitable for ALPs searches.
To accurately calculate the photon survival probability, it is essential to precisely reconstruct the spatial distribution characteristics of the magnetic field within galaxy clusters. This study adopts a systematic magnetic field modeling framework, referencing [51] and its cited series of works [62, 70, 72, 73, 74, 75, 76, 77, 78]. First, a double $\beta$-model is employed to describe the spatial distribution of electron density, expressed as
@@
We investigate the survival probability of photons from the NGC 1275, under different ALP parameters, as illustrated in Figure 1. The left panels show the photon survival probability for different ALP masses, with a constant coupling ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV^{-1}}$). From top to bottom, the ALP mass increases progressively, resulting in the shift of the oscillation energy range: (i) in the top subplot ($m_{a}\ll\omega_{\mathrm{pl}}$), the oscillation range lies to the left of the observation window of Fermi-LAT (100 MeV--500 GeV), but still be in that of VLAST (1 MeV--1 TeV); (ii) in the middle subplot ($m_{a}=1\times 10^{-9}\,\mathrm{eV}$), it falls within both observation windows; (iii) in the bottom subplot ($m_{a}=1\times 10^{-6}\,\mathrm{eV}$), it shifts to the right, outside both observation windows. When the ALP mass is much smaller than the plasma frequency (in the top subplot of left panel), changes in the ALP mass do not affect the trend of photon survival probability according to Equation (2) and Equation (3). The right panel presents the photon survival probability for different couplings while keeping the ALP mass fixed ($m_{a}=1\times 10^{-9}\,\mathrm{eV}$). From top to bottom, the coupling increases gradually. Based on Equation (2) and Equation (3), the characteristic energy decreases continuously. Consequently, the oscillation frequency of the photon survival probability increases, and the oscillation range shifts progressively to the left. These results reveal the distinct dependencies of the ALP conversion process on the ALP mass $m_{a}$ and the coupling $g_{a\gamma}$.
@@ III.1 Fermi-LAT data
Figure 2: The SEDs analysis results for NGC 1275. Top panel: The SEDs of NGC 1275. Black data points show the Fermi-LAT observational results. The blue dotted line represents the best-fit spectrum without an ALP model, while the red solid line shows the best-fit spectrum including an ALP model with $m_{a}=4.72\times 10^{-8}\,\mathrm{eV},g_{a\gamma}=4.89\times 10^{-11}\,\mathrm{GeV^{-1}}$. Bottom panel: $\chi^{2}$ values for each energy bin. Blue solid line with circles indicates $\chi^{2}$ values without the ALP model, and red solid line with triangles correspond to $\chi^{2}$ values including the ALP model with the above specified parameters.
@@
In this work, we utilize nearly 16.5 years of Fermi-LAT P8R3 data [81] collected from October 27, 2008 (MET=246823875), to March 10, 2025 (MET=763293464). Data prior to October 27, 2008 are excluded due to notable background contamination issues at energies above  30 GeV, ensuring the reliability of the analysis  [^3] https://fermi.gsfc.nasa.gov/ssc/data/analysis/LAT_caveats.html.
@@
We select the photon events in a region of interest (ROI) with a radius of $15^{\circ}$ centered on NGC 1275 and in the energy range from 100 MeV to 500 GeV. The “SOURCE” event class and “FRONT+BACK” event type are used for data selection, with the maximum zenith angle set to $90^{\circ}$ to minimize contamination from Earth’s limb. Additionally, the recommended gtmktime filter expression (DATA_QUAL>0) && (LAT_CONFIG==1) is applied to ensure the satellite operates in standard data-taking mode and selects intervals with optimal data quality. Here we use the standard binned likelihood analysis with the Fermi-LAT official software package fermitools (v11r5p3)  [^4] https://github.com/fermi-lat/Fermitools-conda/ with the instrument response function (IRF) P8R3_SOURCE_V3.
@@
With the user-contributed script make4FGLxml.py  [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/user/make4FGLxml.py, the initial model is built as the combination of the Galactic diffuse emission model (gll_iem_v07.fits), the isotropic background spectral template (iso_P8R3_SOURCE_V3_v1.txt), and the incremental version of the fourth full catalog of Fermi-LAT $\gamma$-ray sources (4FGL-DR4) [82]. The Galactic and isotropic background models are available from the Fermi Science Support Center (FSSC)  [^6] https://fermi.gsfc.nasa.gov/ssc/. Spectral parameters for sources within $7^{\circ}$ of NGC 1275 and the normalizations of background models are set as free. We perform the likelihood fit in the energy range (100 MeV--500 GeV) with gtlike tool to obtain the optimized model.
@@
To derive the observed spectral energy distribution (SED) of NGC 1275, we divide the selected events into 37 logarithmically spaced uniform energy bins from 100 MeV to 500 GeV. We fix the spectral indexes of all sources as the values optimized in the global fit, while permitting normalizations to vary freely during the likelihood fit. The test statistic (TS) is defined as minus twice the log-likelihood ratio of the models with and without the target source. Then we repeat the standard binned likelihood analysis in each energy bin to calculate corresponding flux and TS value. For a given energy bin where the TS value is less than 9, we further calculate the 95% confidence level upper limit on the flux using the pyLikelihood UpperLimits tool. The SEDs of NGC 1275 we obtain are shown in Figure 2 as black markers with error bars indicating uncertainties.

### v2

# Constraints on axionlike particles from 16.5 years of Fermi-LAT data and prospects for VLAST

@@ Abstract
Axionlike particles (ALPs), hypothetical particles beyond the Standard Model (SM), are considered as promising dark matter candidates. ALPs can convert into photons and vice versa in a magnetic field via the Primakoff effect, potentially generating detectable oscillation in $\gamma$-ray spectra. This study analyzes 16.5 years of data from the Fermi Large Area Telescope (Fermi-LAT) on NGC 1275, the brightest galaxy in the Perseus cluster, to constrain the ALP parameter space. Our results improve the previous 95% exclusion limits of the photon-ALP coupling $g_{a\gamma}$ by a factor of 2 in the ALP mass range of $4\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$. Moreover, we investigate the projected sensitivity of the future Very Large Area $\gamma$-ray Space Telescope (VLAST) on searching for ALPs. We find that (i) the expected sensitivity on the photon-ALP coupling $g_{a\gamma}$ can be stronger than that from the upcoming International Axion Observatory (IAXO) in the ALP mass range of $2\times 10^{-11}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-7}\,\mathrm{eV}$, with the best sensitivity of $g_{a\gamma}\sim 7\times 10^{-13}\,\mathrm{GeV^{-1}}$ at $m_{a}\sim 2\times 10^{-10}\,\mathrm{eV}$; (ii) VLAST can extend the sensitivity of the ALP masses below $5\times 10^{-12}\,\mathrm{eV}$, where the photon-ALP coupling $g_{a\gamma}\gtrsim 1.5\times 10^{-11}\,\mathrm{GeV^{-1}}$ will be excluded; (iii) the entire parameter space of ALP accounting for TeV transparency can be fully tested. These results demonstrate that VLAST will offer an excellent opportunity for ALPs searches.
@@ I Introduction
The strong CP problem in quantum chromodynamics (QCD) motivated the proposal of the axion, a hypothetical particle beyond the Standard Model (SM) [1, 2, 3, 4], with the mass postulated to be directly proportional to its coupling with photons. More general axionlike particles (ALPs) predicted in many string-theory-motivated extensions of the SM [5, 6, 7], share similarities with axions but are not necessarily tied to the strong CP problem, allowing independent variation of mass and coupling.  [^1] For convenience, we use ALPs to refer to the axions and ALPs in the following context. Besides, the ALPs have been proposed as one of the promising dark matter candidates, whose cosmological abundance can be naturally consistent with observational data via the nonthermal production mechanism [8, 9, 10, 11, 12, 13, 14]. Additionally, ALPs offer a potential solution to various astrophysical anomalies, such as the apparent transparency of TeV photons from extragalactic sources [15, 16, 17, 18]. These motivate the great efforts to detect ALPs through ground-based experiments and astronomical observations  [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41].
@@
Ground-based experiments can be categorized into three main types. The first, light-shining-through-walls experiments, such as the any light particle search at Deutsches Elektronen Synchrotron [42, 43], aim to detect ALPs by shining light through a barrier and observing its conversion into axions. The second type focuses on axion dark matter detection, exemplified by the Axion Dark Matter eXperiment (ADMX) [44, 45, 46, 47, 48, 49, 50, 51], which uses a strong magnetic field to convert dark matter axions into detectable microwave photons. The third type involves solar axion detection, represented by the Axion Solar Telescope (CAST) at CERN [52] and the upcoming International Axion Observatory (IAXO) [53], which aims to greatly enhance sensitivity to solar axions.
Astronomical observations, especially with $\gamma$-ray telescopes, have been believed to be capable of advancing the search for ALPs and effectively narrowing down the viable parameter space for ALPs [54, 55, 56, 57]. Using six years of the Fermi Large Area Telescope (Fermi-LAT) observation on NGC 1275, the couplings above $5\times 10^{-12}\,\mathrm{GeV^{-1}}$ have been excluded for ALP mass $5\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$ [58]. Similarly, the $\gamma$-ray observation for SN 1987A rules out the couplings above $2\times 10^{-13}\,\mathrm{GeV^{-1}}$ for ALP mass $m_{a}\lesssim 1\times 10^{-9}\,\mathrm{eV}$ [59]. Further constraints come from the $\gamma$-ray spectra of extragalactic core-collapse supernovae and bright Fermi flat-spectrum radio quasars (3C 454.3, 3C 279, and CTA 102) [60, 61]. The High Energy Stereoscopic System (HESS) collaboration, using observation from PKS 2155-304, excludes the couplings above $2.1\times 10^{-11}\,\mathrm{GeV^{-1}}$ for ALP mass $1.5\times 10^{-8}\,\mathrm{eV}\lesssim m_{a}\lesssim 6\times 10^{-8}\,\mathrm{eV}$ [62]. Notably, Ref. [63] uses eight years of Fermi-LAT observations on PKS 2155-304 to exclude a previously allowed “hole” parameter region in Ref. [58]. Based on HESS observations, Ref. [64] analyzes the $\gamma$-ray spectra of TeV sources in the Galactic plane, excluding certain parameter space regions and further providing projected sensitivity for the next-generation Cherenkov Telescope Array. Reference [65] combines GeV (Fermi-LAT) and TeV (MAGIC, VERITAS, and HESS) data from three bright supernova remnants to search for ALPs. Reference [66] performs high-resolution polarimetric measurements of radiation near the Galactic Center’s supermassive black hole using a subset of the Event Horizon Telescope array, excluding the couplings above $1\times 10^{-12}\,\mathrm{GeV^{-1}}$ for ALP mass $1\times 10^{-19}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-18}\,\mathrm{eV}$.
@@ II Photon-ALP Oscillation
Table 1: Fundamental model parameters, including those for the double $\beta$ model and the magnetic field model.
@@
Figure 1: Survival probability of photons from NGC 1275 under different ALP parameters based on one realization of the Gaussian turbulent magnetic field, the “Dominguez” EBL model and the “Jansson12” GMF model. In the left panel, with the fixed coupling ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV^{-1}}$), the photon survival probability is presented for different ALP masses (red solid line: $m_{a}\ll\omega_{\mathrm{pl}}$; green solid line: $m_{a}=1\times 10^{-9}\,\mathrm{eV}$; blue solid line: $m_{a}=1\times 10^{-6}\,\mathrm{eV}$). When the ALP mass is significantly smaller than the plasma frequency, its effect on the photon survival probability becomes negligible. In the right panel, with the ALP mass fixed ($m_{a}=1\times 10^{-9}\,\mathrm{eV}$), the photon survival probability is presented for different couplings (red solid line: $g_{a\gamma}=1\times 10^{-13}\,\mathrm{GeV^{-1}}$; green solid line: $g_{a\gamma}=5\times 10^{-12}\,\mathrm{GeV^{-1}}$; blue solid line: $g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV^{-1}}$).
Observational evidence indicates the ubiquitous presence of magnetic fields in Perseus clusters (with the typical strength of $B\sim\mathrm{10\,\mu G}$) and the Milky Way ($B\sim\mathrm{1\,\mu G}$). The $\gamma$-ray photons emitted from NGC 1275 propagate through these magnetic fields and undergo mutual conversions with ALPs via the Primakoff effect. This process, known as photon-ALP oscillation, can introduce irregular oscillatory features into the observed $\gamma$-ray energy spectra of the sources. Such features that are difficult to produce through known astrophysical processes, serve as a smoking-gun signature for the existence of ALPs. Assuming a region of size $\ell$ with a uniform magnetic field, the probability of a photon with energy $E_{\gamma}$ converting into an ALP is given by $P_{\gamma\to a}$ [73, 74, 75, 15, 54, 76, 77, 78, 79, 23, 80], and the photon survival probability $P\left(g_{a\gamma},m_{a},E_{\gamma}\right)$ is
@@
where $\omega_{\mathrm{pl}}$ is the plasma frequency,
@@
This conversion process exhibits strong energy dependence, as detailed in Fig. 1,
(i) $\mathbf{E_{\gamma}\ll E_{c}:}$ the photon-ALP oscillation is weak, with a low conversion probability, and the photon survival probability is close to 1.
(ii) $\mathbf{E_{\gamma}\sim E_{c}:}$ photons and ALPs undergo strong mutual conversion, leading to obvious oscillation.
(iii) $\mathbf{E_{\gamma}\gg E_{c}:}$ the conversion probability is still large, however, the photon survival probability in Eq. (2) mildly varies with $E_{\gamma}$ and becomes gradual.
To distinguish from known astrophysical processes, we focus on searching for the oscillation features with $E_{\gamma}\sim E_{c}$. For typical astrophysical and ALP parameters ($B_{\mathrm{T}}=1\,\mu G,n_{e}=0.01\,\mathrm{cm^{-3}}$ for the galaxy clusters or $n_{e}=0.1\,\mathrm{cm^{-3}}$ for the Milky Way, $m_{a}=1\times 10^{-9}\,\mathrm{eV},g_{a\gamma}=1\times 10^{-10}\,\mathrm{GeV^{-1}}$), the resulting critical energy $E_{c}\approx 250\,\mathrm{MeV}$ falls within the optimal detectable energy ranges of both Fermi-LAT and VLAST, rendering them highly suitable for ALPs searches.
To accurately calculate the photon survival probability, it is essential to precisely reconstruct the spatial distribution characteristics of the magnetic field within galaxy clusters. This study adopts a systematic magnetic field modeling framework, referencing [58] and its cited series of works [70, 79, 81, 82, 83, 84, 85, 86, 87]. First, a double $\beta$ model is employed to describe the spatial distribution of electron density, expressed as
@@
We investigate the survival probability of photons from the NGC 1275, under different ALP parameters, as illustrated in Fig. 1. The left panels show the photon survival probability for different ALP masses, with a constant coupling ($g_{a\gamma}=5\times 10^{-11}\,\mathrm{GeV^{-1}}$). From top to bottom, the ALP mass increases progressively, resulting in the shift of the oscillation energy range: (i) in the top subplot ($m_{a}\ll\omega_{\mathrm{pl}}$), the oscillation range lies to the left of the observation window of Fermi-LAT (100 MeV-500 GeV), but still is in that of VLAST (1 MeV-1 TeV); (ii) in the middle subplot ($m_{a}=1\times 10^{-9}\,\mathrm{eV}$), it falls within both observation windows; (iii) in the bottom subplot ($m_{a}=1\times 10^{-6}\,\mathrm{eV}$), it shifts to the right, outside both observation windows. When the ALP mass is much smaller than the plasma frequency (in the top subplot of the left panel), changes in the ALP mass do not affect the trend of photon survival probability according to Eq. (2) and Eq. (3). The right panel presents the photon survival probability for different couplings while keeping the ALP mass fixed ($m_{a}=1\times 10^{-9}\,\mathrm{eV}$). From top to bottom, the coupling increases gradually. Based on Eq. (2) and Eq. (3), the characteristic energy decreases continuously. Consequently, the oscillation frequency of the photon survival probability increases, and the oscillation range shifts progressively to the left. These results reveal the distinct dependencies of the ALP conversion process on the ALP mass $m_{a}$ and the coupling $g_{a\gamma}$.
@@ III.1 Fermi-LAT data
Figure 2: The SEDs analysis results for NGC 1275. Top panel: The SEDs of NGC 1275. Black data points show the Fermi-LAT observational results. The blue dashdot line represents the best-fit spectrum without an ALP model, while the red solid line shows the best-fit spectrum including an ALP model with $m_{a}=4.72\times 10^{-8}\,\mathrm{eV},g_{a\gamma}=4.89\times 10^{-11}\,\mathrm{GeV^{-1}}$. Bottom panel: $\chi^{2}$ values for each energy bin. Blue solid line with circles indicates $\chi^{2}$ values without the ALP model, and red solid line with triangles correspond to $\chi^{2}$ values including the ALP model with the above specified parameters.
@@
In this work, we utilize nearly 16.5 years of Fermi-LAT P8R3 data [90, 91] collected from October 27, 2008 (MET=246823875), to March 10, 2025 (MET=763293464). Data prior to October 27, 2008 are excluded due to notable background contamination issues at energies above  30 GeV, ensuring the reliability of the analysis.  [^3] https://fermi.gsfc.nasa.gov/ssc/data/analysis/LAT_caveats.html
@@
We select the photon events in a region of interest with a radius of $15^{\circ}$ centered on NGC 1275 and in the energy range from 100 MeV to 500 GeV. The “SOURCE” event class and “FRONT+BACK” event type are used for data selection, with the maximum zenith angle set to $90^{\circ}$ to minimize contamination from Earth’s limb. Additionally, the recommended gtmktime filter expression (DATA_QUAL>0) && (LAT_CONFIG==1) is applied to ensure that the satellite operates in standard data-taking mode and selects intervals with optimal data quality. Here we use the standard binned likelihood analysis with the Fermi-LAT official software package fermitools (v11r5p3)  [^4] https://github.com/fermi-lat/Fermitools-conda/ with the instrument response function P8R3_SOURCE_V3.
@@
With the user-contributed script make4FGLxml.py [^5] https://fermi.gsfc.nasa.gov/ssc/data/analysis/user/make4FGLxml.py the initial model is built as the combination of the Galactic diffuse emission model (gll_iem_v07.fits), the isotropic background spectral template (iso_P8R3_SOURCE_V3_v1.txt), and the incremental version of the fourth full catalog of Fermi-LAT $\gamma$-ray sources (4FGL-DR4) [92]. The Galactic and isotropic background models are available from the Fermi Science Support Center.  [^6] https://fermi.gsfc.nasa.gov/ssc/ Spectral parameters for sources within $7^{\circ}$ of NGC 1275 and the normalizations of background models are set as free. We perform the likelihood fit in the energy range (100 MeV-500 GeV) with the gtlike tool to obtain the optimized model.
@@
To derive the observed spectral energy distribution (SED) of NGC 1275, we divide the selected events into 37 logarithmically spaced uniform energy bins from 100 MeV to 500 GeV. We fix the spectral indexes of all sources as the values optimized in the global fit, while permitting normalizations to vary freely during the likelihood fit. The test statistic (TS) is defined as minus twice the log-likelihood ratio of the models with and without the target source. Then we repeat the standard binned likelihood analysis in each energy bin to calculate corresponding flux and TS value. For a given energy bin where the TS value is less than 9, we further calculate the 95% confidence level upper limit on the flux using the pyLikelihood UpperLimits tool. The SEDs of NGC 1275 we obtain are shown in Fig. 2 as black markers with error bars indicating uncertainties.